The maximum internal compressive stress that the bone can withstand is 1.6 × 108 N/m2. The femur is the largest and longest bone in the human body. The cross-sectional area of ​​the femur is 7.7 × 10-4 m2. Find the maximum intensity of the compressive force that can be applied without breaking the femur. (Answer: F = 123.2 kN)


1
Expert's answer
2021-04-13T06:37:52-0400

"\sigma=F/A\to F=\sigma\cdot A=1.6\cdot10^8\cdot7.7\cdot 10^{-4}="


"=123200\ (N)=123.2\ (kN)" . Answer